OpenStudy (anonymous):

The Grand Canyon is approximately 29 kilometers long. Mariner Valley is a canyon on Mars that is approximately 212 kilometers long. About how many times longer is Mariner Valley than the Grand Canyon? a. 2 b. 2 to the 3rd poer c. 2 to the 21st power d. 2 to the 108 power

OpenStudy (anonymous):

I'll help, but you have to do a part of it first 2^3=what?

OpenStudy (anonymous):

It's just 2*2*2

OpenStudy (anonymous):

2*2*2=8 8*29=232 232 is pretty close to 212 compared to the other answers, so that answer is 2^3

OpenStudy (anonymous):

thanks so much, sorry I didn't reply my computer froze but thank you